OpenStudy (anonymous):

HELP, WILL REWARD!! What is the solution set to the following system? x+y=3 x^2+y^2=9 a.(0,-3)(3,0) b.(0,3)(-3,0) c.(0,3)(3,0) d.(0,-3)(-3,0)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Why??

OpenStudy (anonymous):

if you put them into the equations 0+3=3 and 0^2+3^2=9 and 3+0=3 and 3^2+0^2=9